Herschel Status Report - February 2013
27 March 2013 11:02Report for period 5 to 27 February 2013
Mission operations of the Herschel space observatory continued nominally during the reporting period, with the spacecraft and subsystems all performing as expected.
The spacecraft continues to be in good health and is operating nominally.
At launch, in 2009, Herschel's cryostat was filled with over 2300 litres of superfluid liquid helium, weighing about 335 kg. The helium is steadily used by the spacecraft's active cooling system to cool the payload, so the amount of remaining coolant directly determines Herschel's lifetime for scientific observations. Best estimates for the date of the exhaustion of the liquid helium (LHe) continue to fall in the second-half of March this year.
After Herschel has completed its operational lifetime for scientific observations, the spacecraft will leave its current orbit around L2 and will be injected into a no-return heliocentric orbit. In this context "no-return" means not returning to the potential well of the Earth-Moon system for at least 300 years. The window for the manoeuvre required to put the spacecraft in its final heliocentric orbit opens on 5 May 2013.
A long period free of big contingencies for any of the three instruments, PACS, SPIRE and HIFI, ended with four anomalies in the last two weeks covered by this report. On 12 February, SPIRE was switched off by autonomous on-board control procedure (OBCP). About thirteen hours later, HIFI too was switched off by autonomous OBCP. SPIRE could be quickly restored to nominal operation. HIFI's recovery to nominal operation took longer and was completed on 21 February.
While the HIFI recovery was progressing PACS was found, through a routine quality check of science data, to have one of its two photometer red bolometer arrays saturated. Later, on 23 February, PACS was found switched off by autonomous OBCP. The PACS spectrometer was fully recovered at the following pass and is again nominal. The PACS photometer is operational again apart from a, most-likely irrecoverable, failure of the one red bolometer array.
The series of apparently unrelated anomalies resulted in extensive round-the-clock recovery actions, involving instrument manual operations and replanning of science schedules, also to minimise the science time lost.
Ground segment operations have been nominal and 100% of the data continues to be recovered. As of 22 February 2013, the approximate completion of the different programme parts was:
A bulk reprocessing of all the science data in the HSA, using the then-latest release of the data processing software (version 9), was started in November 2012 and has been completed in this reporting period.